OpenStudy (anonymous):

Find the y-coordinate of the vertex for the following function: g(x) = x2 - 3x + 2 y= its asking for a simplified fraction?

OpenStudy (mertsj):

you might want to put it into vertex form by completing the square. \[x^2 -3x +(\frac{3}{2})^{2 }+2-(\frac{3}{2})^{2}=\] \[(x-\frac{3}{2})^{2}+(-\frac{1}{4})\]

OpenStudy (mertsj):

so we see that the vertex is( 3/2,-14)

OpenStudy (anonymous):

so the fraction would be 3/2?

OpenStudy (amistre64):

when x=3/2; what does y=?

OpenStudy (amistre64):

-1/4 ; he dropped the bar

OpenStudy (anonymous):

oh, lol thank you!